# LLM-Driven Corrective Robot Operation Code Generation with Static Text-Based Simulation

### v2

@@ I INTRODUCTION
Robots are being increasingly deployed to execute tasks based on human instructions. However, designing a robot that has intelligence to perform complex tasks reliably remains challenging, as it demands both robust instruction interpretation and executing tasks with robust reasoning. Recent advances in LLMs [GPT4, Gemini, deepseek] have demonstrated remarkable proficiency in robotic areas such as control [Real], planning [planning], and navigation [L3MVN]. Their strong context understanding and generation capabilities enable the robot to comprehend human instructions and generate corresponding robot operation code [ChatGPTRobotics, TypeFly, CodeasPolicies, GSCE], thereby greatly simplifying the process of robot programming.
Unlike text generation applications, where semantic-level accuracy is sufficient, executing logically inconsistent or syntactically incorrect code on a robot can lead to unexpected outcomes and even unsafe robot behaviors, such as UAV crashes. To improve the reliability of LLMs’ robot operation code generation, recent studies have incorporated a corrective process that leverages the observation of executing LLM-generated code in a physical experiment or simulator to identify issues and perform refinement [AutoTAMP, InteractivePlanning, CLGSCE]. While these corrective designs have shown their effectiveness in enhancing the reliability of LLM-generated operation code, they still face challenges from different aspects. Specifically, the configuration of a physical or simulation environment for robot code execution requires specialized expertise to support the tasks (e.g., designing scenes and modeling a robot’s replica), especially for custom-built robots. For example, the code execution in ref [CLGSCE] requires a specific design of mapping that transforms numerical state representation to the semantic description, which needs to be customized case-by-case. In addition, existing corrective designs require multiple rounds of interactive refinements, which can lead to a long execution time in the experiment or simulation when the task is complicated or involves time-consuming operations (e.g., monitor the area for 15 minutes). Therefore, removing the dependence on dynamic execution in a physical experiment or simulator while still maintaining the corrective feedback and refinement features for LLM-driven robot operation code generation becomes a challenging gap to close.
@@
We extensively evaluated our proposed framework for the operation code generation of UAV tasks with various levels of task complexities. Our results show that our static text-based simulation achieves over $97.5\%$ of simulation accuracy compared with the widely adopted UAV simulators, i.e., AirSim [AirSim] and PX4-Gazebo [PX4]. In addition, our corrective code generation framework delivers comparable robot execution performance as the state-of-the-art (SOTA) method relying on dynamic code execution in physical experiment or simulator, with an $85\%+$ success rate and $96.9\%+$ completeness on different UAV systems, i.e., $96.9\%+$ of required actions in all evaluated tasks are completed correctly and $85\%+$ of tasks are entirely completed without any error. To demonstrate the adaptability of our framework and its performance on real-world robot deployment, we also evaluated it for the operation code generation for UAVs and ground robots. Our experiment results show that our framework can also achieve high success rates ($87.5\%+$) and completeness ($96.9\%+$).
@@ II-A LLM-Driven Corrective Robot Code Generation
Using LLMs to generate operation codes for robot tasks has become a prevalent trend in recent research. By configuring LLMs with appropriate system prompting techniques, existing research has shown that LLMs have the potential to generate robot operation codes  [ChatGPTRobotics, TypeFly, CodeasPolicies, GSCE]. To further enhance the reliability of LLMs’ output and support more complicated robot tasks, recent efforts have adopted corrective code generation such that the errors or mismatches are iteratively detected and corrected; therefore, the robot could perform the desired task accurately when executing the final code [CoPAL, CAPE, trust]. However, the physical execution during the correction process may increase the hardware cost and raise safety risks, as executing code with errors could cause irreparable damage to robots, such as UAV crashes. More recent studies proposed a simulation-based pipeline that eliminates the potential risks during physical execution [CLGSCE, AutoTAMP, hu2024robo]. For example, [CLGSCE] leverages the AirSim [AirSim] simulator to iteratively correct the UAV operation code until the code is ready for deployment. The authors also develop semantic observations rather than numerical representation [InteractivePlanning] to describe the UAV trajectory to further improve performance. Robo-Instruct [hu2024robo] fine-tuning LLM that checks LLM-generated robot programs with a simulator and revises them until correct.
@@ II-B LLM-based Simulation
Recent advances in LLMs [GPT4, Gemini, deepseek] have demonstrated their exceptional capabilities in world modeling. Previous studies have explored the integration of LLMs into agent-based modeling and simulation within the social [LLM-simulator-survey] and planning [yang2024evaluating] domains. For example, BeSimulator [besimulator] built an LLM-powered framework towards behavior simulation on the behavior tree. However, studies on text game show that current LLMs are not yet able to reliably act as text world simulators as they are likely to make errors when arithmetic, common-sense, or scientific knowledge is needed [LLM-simulator]. This limitation highlights the need for further research to examine the use of LLMs for the robot code simulation domain and develop methods to enhance the ability of LLMs to perform accurate and reliable simulations of robot operation code.
@@ II-C Prompt Engineering
Prompt engineering can effectively communicate and interact with LLM-driven tools [PromptEngineering]. Recent studies have increasingly used prompt engineering to improve the reliability of LLM-driven robotic systems. For example, PromptBook proposes a prompt framework as a system prompt to enhance the code generation [PromptBook]. On the one hand, prompt engineering leverages in-context learning [URIAL] and few-shot learning [llm-fewshot] to enable LLM to learn knowledge within the given context and identify patterns from a limited set of examples. These techniques facilitate the generation of code that adheres to robot policy and how to ground task descriptions from a few examples [CodeasPolicies]. On the other hand, CoT [CoT] prompts the LLM to articulate intermediate inference steps, which is valuable for robotic tasks that require sequential, stepwise decision-making. CoT encourages the production of code that aligns with each stage of the intended action plan. Previous studies have embedded CoT within the examples to guide LLMs through reasoning step-by-step  [ISR-LLM, PromptBook]. In this study, we utilize prompt engineering strategies to facilitate our text-based simulation.
@@ III-B Code Generation
The code generation in our method is achieved by configuring an LLM agent (code generator) with a robot operation-related system prompt. We adopt the strategies in the GSCE framework [GSCE], which enhance the reasoning capabilities of LLMs in generating UAV operation code that aligns with task instructions. Additionally, the code generator regenerates the code that resolves the mismatches according to the feedback from the evaluator in section III-D. In detail, given a task description or evaluation feedback, the code generator produces a robot operation code that aims to accomplish the task or resolve the mismatches.
@@ III-C Static Text-Based Simulation using LLM
  • Policies: Instruct the LLM with the code execution polices where the LLM lacks prior knowledge. The policies clarify the APIs usage, state transition rules, and environmental settings.
@@ III-D Evaluation
In evaluation, the evaluator identifies the mismatches (if any) between the observation $O$ and the task description, and then provides feedback specifying the mismatched actions. To ensure evaluation accuracy, we adopted the evaluator design in [CLGSCE], which has demonstrated effectiveness in identifying the deviations between the UAV’s trajectory and task description. The feedback from the evaluation provides the code generator with a clearer understanding of the objectives implied by the task description, thereby steering the generation of corrected code to better align with the task.
@@ IV-A1 Experimental Environment
We implement our proposed method and comparison method using OpenAI “o3-mini” (o3-mini-2025-01-31) [o3mini] and “o4-mini” (o4-mini-2025-04-16) [o4mini] as the foundational LLMs. To measure the performance of the methods, the experiment is conducted on a quadcopter on both the simple_flight flight controller in AirSim [AirSim] and the PX4 flight controller [PX4] in Gazebo [gazebo]  [^2] Simulator configurations and setups are provided in Appendix B. During the experiment, UAV state information was accessible from the simulators and utilized for performance measurement purposes. Furthermore, all experiments are averaged over three repetitions to mitigate the randomness of LLM generation [LLMrandomness].
@@ IV-A2 Task Dataset
For the experiments, we adopt the Advanced task set from [CLGSCE] as the benchmark dataset to measure the performance. The Advanced task set contains 20 UAV operation tasks with varying levels of complexity, each involving 6-19 actions to reach the goal state. The tasks are designed to emulate real-world UAV operation scenarios that require complex reasoning about the UAV’s state in the world environment, such as flying complex geometric patterns with scenario requirements. To ensure the authenticity of the result, all tasks are manually validated in both AirSim and Gazebo to avoid potential simulation-induced errors that could affect the experiment result.
@@ IV-A3 Compared Method
  • Direct Analysis (Direct): Uses the semantic checker from [AutoTAMP], where an LLM agent directly analyzes whether the generated code aligns with the task and provides feedback for code correction.
  • Simulator-based (Numerical): Dynamically executes the generated code in a simulator to obtain numerical state observations, which are then evaluated to provide feedback for code correction [InteractivePlanning].
  • Simulator-based (Semantic): A SOTA method extends the Numerical method [InteractivePlanning] by transforming numerical state observations produced from the simulator into semantic trajectory descriptions [CLGSCE] to improve the code correction efficiency. During the experiments, we modified the transformation algorithm to accommodate different robots and simulators.
@@ IV-B1 Evaluation Metrics
Following [CLGSCE], we evaluate performance using Completeness and Success Rate (SR).
@@ IV-C1 Overall Result
In particular, our method achieves performance comparable to the SOTA Semantic method [CLGSCE] without requiring dynamic code execution in simulators that are explicitly designed to support robot simulations. This highlights that the proposed LLM-simulator can reliably conduct static text-based simulation of robot code to support corrective code generation. Furthermore, our method outperforms the Numerical method [InteractivePlanning], indicating that the semantic observations generated by our LLM-simulator capture richer semantics about the robot’s trajectory dynamics than the numerical representations. This enables an equally effective correction process as the SOTA Semantic method [CLGSCE], but without the need for customizing algorithms for different robot configurations to transform numerical states into semantic descriptions. In contrast, the Direct method [AutoTAMP] yields unreliable performance, proving the limitations of directly configuring LLMs to analyze code and highlighting the necessity to strengthen LLM’s capabilities for LLM-based simulation.
@@
Direct [AutoTAMP] $43.3\%$ $70.9\%$ $33.3\%$ $57.6\%$
Numerical [InteractivePlanning] $73.3\%$ $92.4\%$ $81.7\%$ $96.1\%$
Semantic [CLGSCE] $\mathbf{85.0}\%$ $\mathbf{98.5}\%$ $88.3\%$ $98.1\%$
@@
Direct [AutoTAMP] $55.0\%$ $77.5\%$ $25.0\%$ $50.9\%$
Numerical [InteractivePlanning] $83.3\%$ $97.1\%$ $86.7\%$ $96.9\%$
Semantic [CLGSCE] $\mathbf{88.3}\%$ $\mathbf{98.3}\%$ $\mathbf{93.3}\%$ $\mathbf{98.6}\%$
@@ IV-C2 Text-Based Simulation Accuracy
To further evaluate the reliability of our text-based simulation, we compare the accuracy of the trajectory observations generated by the LLM-simulator against those obtained using the Semantic method [CLGSCE] in the dynamic simulator. For each task in the Advanced task set, we construct a corresponding ground truth code ($C^{\text{correct}}$) that implements the task. The LLM-simulator is then used to statically simulate the execution of each $C^{\text{correct}}_{i}$ and produce trajectory observations. The accuracy of the simulation is computed as $TP/N\times 100\%$, where $TP$ denotes the number of observations that accurately capture the robot trajectory dynamics, and $N$ is the total number of simulated code in $C^{\text{correct}}$.
@@
TABLE III: Text-Base Simulation Observation Accuracy
@@ IV-C3 Text-Based Simulation Evaluation Accuracy
We further analyze whether the observations generated by the LLM-simulator can effectively support the evaluation process. Specifically, we compare the evaluation accuracy when using observations generated by our LLM-simulator against those produced from the Semantic method [CLGSCE]. This experiment leverages the ground-truth code set $C^{\text{correct}}$ in section IV-C2, along with an additional set $C^{\text{incorrect}}$ that contains at least one error action in code for each task. The evaluation accuracy is computed as $(TP+TN)/N\times 100\%$, where $TP$ denotes the number of correct evaluations on $C^{\text{correct}}$ and $C^{\text{incorrect}}$ (i.e., correctly identifying whether the trajectory matches with the task and providing a detailed explanation if mismatches are identified), and $N$ is the total number of evaluated code scripts.
As presented in Table IV, our method achieves over $90\%$ accuracy on “o3-mini” and $91.7\%$ accuracy on “o4-mini”. The results are comparable to the Semantic method [CLGSCE] without the need for modifying algorithms for each robot and simulator to transform numerical states into semantic descriptions. The results also demonstrate that the observations from the LLM-simulator are sufficient to support effective evaluation as the SOTA Semantic method [CLGSCE] and are adaptable to different robot systems.
However, due to the non-determinism of LLM generation [Non-determinism], identical actions may yield semantically equivalent but syntactically different observations. For example, the action “fly 5 meters south” from the text-based simulation may be produced as “followed by 5 meters back (returning in the north-south direction)”. While such descriptions remain interpretable to humans, their implicit ambiguity can cause misinterpretations for LLMs [ambiguity-misinterpret]. Consequently, observations generated by the Semantic method [CLGSCE] exhibit slightly higher evaluation accuracy, as they are deterministic and free from linguistic variations.

### v3

@@ I INTRODUCTION
Robots are being increasingly deployed to execute tasks based on human instructions. However, designing a robot that has intelligence to perform complex tasks reliably remains challenging, as it demands both robust instruction interpretation and executing tasks with robust reasoning. Recent advances in LLMs [1, 30, 16] have demonstrated remarkable proficiency in robotic areas such as control [29], planning [21], and navigation [38]. Their strong context understanding and generation capabilities enable the robot to comprehend human instructions and generate corresponding robot operation code [31, 5, 14, 34], thereby greatly simplifying the process of robot programming.
Unlike text generation applications, where semantic-level accuracy is sufficient, executing logically inconsistent or syntactically incorrect code on a robot can lead to unexpected outcomes and even unsafe robot behaviors, such as UAV crashes. To improve the reliability of LLMs’ robot operation code generation, recent studies have incorporated a corrective process that leverages the observation of executing LLM-generated code in a physical experiment or simulator to identify issues and perform refinement [6, 28, 35]. While these corrective designs have shown their effectiveness in enhancing the reliability of LLM-generated operation code, they still face challenges from different aspects. Specifically, the configuration of a physical or simulation environment for robot code execution requires specialized expertise to support the tasks (e.g., designing scenes and modeling a robot’s replica), especially for custom-built robots. For example, the code execution in ref [35] requires a specific design of mapping that transforms numerical state representation to the semantic description, which needs to be customized case-by-case. In addition, existing corrective designs require multiple rounds of interactive refinements, which can lead to a long execution time in the experiment or simulation when the task is complicated or involves time-consuming operations (e.g., monitor the area for 15 minutes). Therefore, removing the dependence on dynamic execution in a physical experiment or simulator while still maintaining the corrective feedback and refinement features for LLM-driven robot operation code generation becomes a challenging gap to close.
@@
We extensively evaluated our proposed framework for the operation code generation of UAV tasks with various levels of task complexities. Our results show that our static text-based simulation achieves over $97.5\%$ of simulation accuracy compared with the widely adopted UAV simulators, i.e., AirSim [27] and PX4-Gazebo [19]. In addition, our corrective code generation framework delivers comparable robot execution performance as the state-of-the-art (SOTA) method relying on dynamic code execution in physical experiment or simulator, with an $85\%+$ success rate and $96.9\%+$ completeness on different UAV systems, i.e., $96.9\%+$ of required actions in all evaluated tasks are completed correctly and $85\%+$ of tasks are entirely completed without any error. To demonstrate the adaptability of our framework and its performance on real-world robot deployment, we also evaluated it for the operation code generation for UAVs and ground robots. Our experiment results show that our framework can also achieve high success rates ($87.5\%+$) and completeness ($96.9\%+$).
@@ II-A LLM-Driven Corrective Robot Code Generation
Using LLMs to generate operation codes for robot tasks has become a prevalent trend in recent research. By configuring LLMs with appropriate system prompting techniques, existing research has shown that LLMs have the potential to generate robot operation codes  [31, 5, 14, 34]. To further enhance the reliability of LLMs’ output and support more complicated robot tasks, recent efforts have adopted corrective code generation such that the errors or mismatches are iteratively detected and corrected; therefore, the robot could perform the desired task accurately when executing the final code [11, 26, 7]. However, the physical execution during the correction process may increase the hardware cost and raise safety risks, as executing code with errors could cause irreparable damage to robots, such as UAV crashes. More recent studies proposed a simulation-based pipeline that eliminates the potential risks during physical execution [35, 6, 10]. For example, [35] leverages the AirSim [27] simulator to iteratively correct the UAV operation code until the code is ready for deployment. The authors also develop semantic observations rather than numerical representation [28] to describe the UAV trajectory to further improve performance. Robo-Instruct [10] fine-tuning LLM that checks LLM-generated robot programs with a simulator and revises them until correct.
@@ II-B LLM-based Simulation
Recent advances in LLMs [1, 30, 16] have demonstrated their exceptional capabilities in world modeling. Previous studies have explored the integration of LLMs into agent-based modeling and simulation within the social [9] and planning [37] domains. For example, BeSimulator [32] built an LLM-powered framework towards behavior simulation on the behavior tree. However, studies on text game show that current LLMs are not yet able to reliably act as text world simulators as they are likely to make errors when arithmetic, common-sense, or scientific knowledge is needed [33]. This limitation highlights the need for further research to examine the use of LLMs for the robot code simulation domain and develop methods to enhance the ability of LLMs to perform accurate and reliable simulations of robot operation code.
@@ II-C Prompt Engineering
Prompt engineering can effectively communicate and interact with LLM-driven tools [17]. Recent studies have increasingly used prompt engineering to improve the reliability of LLM-driven robotic systems. For example, PromptBook proposes a prompt framework as a system prompt to enhance the code generation [2]. On the one hand, prompt engineering leverages in-context learning [15] and few-shot learning [4] to enable LLM to learn knowledge within the given context and identify patterns from a limited set of examples. These techniques facilitate the generation of code that adheres to robot policy and how to ground task descriptions from a few examples [14]. On the other hand, CoT [36] prompts the LLM to articulate intermediate inference steps, which is valuable for robotic tasks that require sequential, stepwise decision-making. CoT encourages the production of code that aligns with each stage of the intended action plan. Previous studies have embedded CoT within the examples to guide LLMs through reasoning step-by-step  [39, 2]. In this study, we utilize prompt engineering strategies to facilitate our text-based simulation.
@@ III-B Code Generation
The code generation in our method is achieved by configuring an LLM agent (code generator) with a robot operation-related system prompt. We adopt the strategies in the GSCE framework [34], which enhance the reasoning capabilities of LLMs in generating UAV operation code that aligns with task instructions. Additionally, the code generator regenerates the code that resolves the mismatches according to the feedback from the evaluator in section III-D. In detail, given a task description or evaluation feedback, the code generator produces a robot operation code that aims to accomplish the task or resolve the mismatches.
@@ III-C Static Text-Based Simulation using LLM
  • Policies: Instruct the LLM with the code execution policies where the LLM lacks prior knowledge. The policies clarify the APIs usage, state transition rules, and environmental settings.
@@ III-D Evaluation
In evaluation, the evaluator identifies the mismatches (if any) between the observation $O$ and the task description, and then provides feedback specifying the mismatched actions. To ensure evaluation accuracy, we adopted the evaluator design in [35], which has demonstrated effectiveness in identifying the deviations between the UAV’s trajectory and task description. The feedback from the evaluation provides the code generator with a clearer understanding of the objectives implied by the task description, thereby steering the generation of corrected code to better align with the task.
@@ IV-A1 Experimental Environment
We implement our proposed method and comparison method using OpenAI “o3-mini” (o3-mini-2025-01-31) [22] and “o4-mini” (o4-mini-2025-04-16) [23] as the foundational LLMs. To measure the performance of the methods, the experiment is conducted on a quadcopter on both the “simple_flight” flight controller in AirSim [27] and the PX4 flight controller [19] in Gazebo [13]  [^2] Simulator configurations and setups are provided in Appendix B. During the experiment, UAV state information was accessible from the simulators and utilized for performance measurement purposes. Furthermore, all experiments are averaged over three repetitions to mitigate the randomness of LLM generation [3].
@@ IV-A2 Task Dataset
For the experiments, we adopt the Advanced task set from [35] as the benchmark dataset to measure the performance. The Advanced task set contains 20 UAV operation tasks with varying levels of complexity, each involving 6-19 actions to reach the goal state. The tasks are designed to emulate real-world UAV operation scenarios that require complex reasoning about the UAV’s state in the world environment, such as flying complex geometric patterns with scenario requirements. To ensure the authenticity of the result, all tasks are manually validated in both AirSim and Gazebo to avoid potential simulation-induced errors that could affect the experiment result.
@@ IV-A3 Compared Method
  • Direct Analysis (Direct): Uses the semantic checker from [6], where an LLM agent directly analyzes whether the generated code aligns with the task and provides feedback for code correction.
  • Simulator-based (Numerical): Dynamically executes the generated code in a simulator to obtain numerical state observations, which are then evaluated to provide feedback for code correction [28].
  • Simulator-based (Semantic): A SOTA method extends the Numerical method [28] by transforming numerical state observations produced from the simulator into semantic trajectory descriptions [35] to improve the code correction efficiency. During the experiments, we modified the transformation algorithm to accommodate different robots and simulators.
@@ IV-B1 Evaluation Metrics
Following [35], we evaluate performance using Completeness and Success Rate (SR).
@@ IV-C1 Overall Result
In particular, our method achieves performance comparable to the SOTA Semantic method [35] without requiring dynamic code execution in simulators that are explicitly designed to support robot simulations. This highlights that the proposed LLM-simulator can reliably conduct static text-based simulation of robot code to support corrective code generation. Furthermore, our method outperforms the Numerical method [28], indicating that the semantic observations generated by our LLM-simulator capture richer semantics about the robot’s trajectory dynamics than the numerical representations. This enables an equally effective correction process as the SOTA Semantic method [35], but without the need for customizing algorithms for different robot configurations to transform numerical states into semantic descriptions. In contrast, the Direct method [6] yields unreliable performance, proving the limitations of directly configuring LLMs to analyze code and highlighting the necessity to strengthen LLM’s capabilities for LLM-based simulation.
@@
Direct [6] $43.3\%$ $70.9\%$ $33.3\%$ $57.6\%$
Numerical [28] $73.3\%$ $92.4\%$ $81.7\%$ $96.1\%$
Semantic [35] $\mathbf{85.0}\%$ $\mathbf{98.5}\%$ $88.3\%$ $98.1\%$
@@
Direct [6] $55.0\%$ $77.5\%$ $25.0\%$ $50.9\%$
Numerical [28] $83.3\%$ $97.1\%$ $86.7\%$ $96.9\%$
Semantic [35] $\mathbf{88.3}\%$ $\mathbf{98.3}\%$ $\mathbf{93.3}\%$ $\mathbf{98.6}\%$
@@ IV-C2 Text-Based Simulation Accuracy
To further evaluate the reliability of our text-based simulation, we compare the accuracy of the trajectory observations generated by the LLM-simulator against those obtained using the Semantic method [35] in the dynamic simulator. For each task in the Advanced task set, we construct a corresponding ground truth code ($C^{\text{correct}}$) that implements the task. The LLM-simulator is then used to statically simulate the execution of each $C^{\text{correct}}_{i}$ and produce trajectory observations. The accuracy of the simulation is computed as $TP/N\times 100\%$, where $TP$ denotes the number of observations that accurately capture the robot trajectory dynamics, and $N$ is the total number of simulated code in $C^{\text{correct}}$.
@@
TABLE III: Text-Based Simulation Observation Accuracy
@@ IV-C3 Text-Based Simulation Evaluation Accuracy
We further analyze whether the observations generated by the LLM-simulator can effectively support the evaluation process. Specifically, we compare the evaluation accuracy when using observations generated by our LLM-simulator against those produced from the Semantic method [35]. This experiment leverages the ground-truth code set $C^{\text{correct}}$ in section IV-C2, along with an additional set $C^{\text{incorrect}}$ that contains at least one error action in code for each task. The evaluation accuracy is computed as $(TP+TN)/N\times 100\%$, where $TP$ denotes the number of correct evaluations on $C^{\text{correct}}$ and $C^{\text{incorrect}}$ (i.e., correctly identifying whether the trajectory matches with the task and providing a detailed explanation if mismatches are identified), and $N$ is the total number of evaluated code scripts.
As presented in Table IV, our method achieves over $90\%$ accuracy on “o3-mini” and $91.7\%$ accuracy on “o4-mini”. The results are comparable to the Semantic method [35] without the need for modifying algorithms for each robot and simulator to transform numerical states into semantic descriptions. The results also demonstrate that the observations from the LLM-simulator are sufficient to support effective evaluation as the SOTA Semantic method [35] and are adaptable to different robot systems.
However, due to the non-determinism of LLM generation [24], identical actions may yield semantically equivalent but syntactically different observations. For example, the action “fly 5 meters south” from the text-based simulation may be produced as “followed by 5 meters back (returning in the north-south direction)”. While such descriptions remain interpretable to humans, their implicit ambiguity can cause misinterpretations for LLMs [12]. Consequently, observations generated by the Semantic method [35] exhibit slightly higher evaluation accuracy, as they are deterministic and free from linguistic variations.